OpenStudy (vespersaviour):

Which expressions could complete this equation so that it has one solution? Choose exactly two answers that are correct.

OpenStudy (vespersaviour):

Attachment 1 is the equation Attachment 2 is the answers

OpenStudy (vespersaviour):

@ILovePuppiesLol can you help?

ILovePuppiesLol (ilovepuppieslol):

i don't know D:

OpenStudy (vespersaviour):

The question \[4\left( 4x-3 \right)-6x\] The answers \[12x-12-2x\] \[9x-12\] \[4x+2\left( 3x-6 \right)\] \[3x+3\left( 2x-4 \right)\]

Directrix (directrix):

The first thing to do is to simplify the given expression: 4(4x - 3) - 6x = ?

OpenStudy (vespersaviour):

Would it be 4(10x - 3) = ?

Directrix (directrix):

Use the distributive property to expand this: 4(4x - 3)

Directrix (directrix):

>>4(10x - 3) = No. This is not correct.

Directrix (directrix):

Expand this: 4*(4x - 3)